Question 1
Howard Bowen is a large-scale cotton farmer. The land and machinery he owns has a
current market value of $4 million. Bowen owes his local bank $3 million. Last year,
Bowen sold $5 million worth of cotton. His variable operating costs were $4.5 million;
accounting depreciation was $40,000, although the actual decline in value of Bowen’s
machinery was $60,000 last year. Bowen paid himself a salary of $50,000, which is not
considered part of his variable operating costs. Interest on his bank loan was $400,000.
If Bowen worked for another farmer or a local manufacturer, his annual income would be
about $30,000. Bowen can invest any funds that would be derived, if the farm sold, to
earn 10 percent annually. (Ignore taxes.)
a. Compute and interpret Bowen’s accounting profits.
b. Compute and interpret Bowen’s economic profits.
1
Expert's answer
2013-02-06T11:39:40-0500
The accounting profit is the difference between total revenue and total cost excluding the economic cost (opportunity cost) of owner-supplied resources such as time and capital. On the other hand, In the economic cost, we include the opportunity cost in our calculations.

Accounting profit = 5,000,000 - 4,500,000 - 40,000 - 400,000 - 50,000 = $10,000
Economic profit = 5,000,000 - 4,500,000 - 60,000 - 400,000 - 30,000 - 4,000,000*10% = -$390,000

Howard Bowen is a large-scale cotton farmer. The land and machinery he owns has a current market value of $4 million. Bowen owes his local bank $3 million. Last year, Bowen sold $5 million worth of cotton. His variable operating costs were $4.5 million; accounting depreciation was $40,000, although the actual decline in value of Bowen’s machinery was $60,000 last year. Bowen paid himself a salary of $50,000, which is not considered part of his variable operating costs. Interest on his bank loan was $400,000. If Bowen worked for another farmer or a local manufacturer, his annual income would be about $30,000. Bowen can invest any funds that would be derived, if the farm sold, to earn 10 percent annually. (Ignore taxes.)